# PGD-based optimization of 3D bobsleigh track centerlines from 2D centerlines for simulation applications

### v1

@@ Abstract
The centerline of a bobsleigh track defines its geometry and is essential for simulation modeling. To reduce bBobsleigh training costs, leveraging the centerline of the bobsleigh track to construct a virtual environment that closely replicates real competitive settings presents a promising solution. However, publicly available centerline data are typically limited and it is imprecise to construct a training system solely based on 2-dimensional (2D) centerline. To address this practical issue, this paper proposes a method for generating a 3-dimensional (3D) track centerline based on 2D centerline data. Incorporating international track design regulations, the method formulates an optimization problem that considers total track length, height difference, slope constraints, and geometric continuity. A Projected Gradient Descent (PGD) algorithm is used to solve the optimization problem. The generated 3D centerlines are compared with real track data, and the results show that the method can reproduce realistic centerline trends from original or scaled 2D data. For the selected track segment, the relative errors in total length, height difference, and average slope are within 1.7%, 3.2% and 4.1%, respectively, for real 2D data and within 1.1%, 3.5% and 4.3% respectively for scaled data. All slope values remain within the allowable limits. Moreover, by adjusting the segmentation or modifying the weight of height difference in the cost function, various centerline styles applicable to different competitions can be generated. Under different segmentation and weight factors, the maximum errors reach up to 4.4%, 4.8%, and 9.8%, and 4.4%, 4.8%, and 10.0%, respectively. The proposed method provides a flexible and efficient tool for supporting bobsleigh track centerline design.
@@
Bobsleigh, track centerline generation, international Olympic bobsleigh competition rules, projected gradient descent
@@ 1 Introduction
Since bobsleigh was formally recognized as a Winter Olympics event in 1924, it has attracted widespread attention from participating countries bi:bu4 . Bobsleighs are unpowered vehicles that rely entirely on the initial push from athletes and the conversion of gravitational potential energy into kinetic energy as they descend the track bi:1 ; 3 ; 5 . The competition is timed to an accuracy of milliseconds, and the final times often differ by only a few tenths of a second4 . To minimize energy loss, athletes typically avoid using the brake unless necessary before crossing the finish line. The dynamics of the bobsleigh and the driver’s control skill are key factors influencing performance bi:bu1 ; r2 .
@@
However, due to restrictions from official governing bodies and the high cost of maintaining bobsleigh tracks7 , even countries with access to tracks often do not allow athletes to train on them regularly. Typically, limited test runs are permitted only shortly before official competitions bi:2 ; 1 ; bi:bu2 ; 6 . Consequently, simulation-based training and performance analysis have become vital tools for optimizing bobsleigh design and improving athlete technique. Researchers have developed virtual tracks, bobsleigh-blade/ice interaction models, multibody dynamics models of bobsleighs, and driver models. These simulation frameworks are increasingly being used to support bobsleigh structural design and help athletes devise optimal racing strategies. And there are limited studies on optimizing the centerline of bobsleigh tracks.
@@
Early bobsleigh simulation research was significantly constrained by the limited computational capabilities of CPUs at the time, often relying on simplified point-mass models moving on smooth hyperbolic surfaces bi:3 ; bi:4 . The study bi:4 developed a partial surface model of a bobsleigh track, treating the bobsleigh as a point mass to analyze its dynamic response under external forces. However, the authors also acknowledged that due to the lack of available data, virtual training environments differ significantly from the real settings, limiting the variety of technical strategies athletes can efficiently practice. Based on this point-mass framework and real data from the Lillehammer track in Norway, The study bi:5 further designed an optimal controller to simulate bobsleigh driving behavior, aiding in athlete training. To better capture the complex nonlinear dynamics of bobsleigh motion on narrow, twisted, banked, and icy tracks, a geometric model is proposed bi:6 , which represents the tracks as a sequence of discrete points arranged along a 3D spatial curve. Moreover, a comprehensive simulation framework incorporating a blade–ice contact friction model is designed and validated on the Cesana and Whistler tracks, offering valuable guidance for track design and athlete training. The study bi:8 further validates its agreement between the simulated and experimental results on the Cesana track. However, the study did not involve real-time simulation, and the applicability of the proposed model to other official tracks worldwide was not examined. Subsequently, the geometric model is further extended into a multibody dynamics model in bi:2 , where more detailed dynamic characteristics are considered, resulting in better driving fidelity. Recent advances in bobsleigh simulation have focused on improving blade–ice friction modeling and virtual training environments bi:bu3 ; bi:7 . The study bi:10 developed longitudinal and lateral friction models through experiments, FEM and nonlinear regression, validating their effectiveness in simulators to evaluate driving styles. Some studies further integrate steering and visual feedback or design immersive systems for training and tourism bi:3 ; bi:2 ; bi:10 ; bi:11 . The study bi:11 introduced a VR hammock system to mimic sliding sensations, while the studies bi:12 ; bi:13 proposed automated methods for generating 3D track surfaces.
@@
However, most of the aforementioned studies focus on the simulation of only one or two tracks, and the conclusions drawn lack generalizability. Many researchers have acknowledged the necessity of validating their proposed models against simulation results from other tracks in order to derive more universally applicable insights for athlete training. Currently, track data are primarily obtained through a combination of field measurements and industrial software modeling, which involves significant labor and material costs. Moreover, most countries do not publicly disclose the official data of their bobsleigh tracks. Therefore, acquiring a greater variety of available official track data worldwide has become an urgent need for enhancing both the diversity of research findings in simulation analysis and the general applicability of virtual training outcomes for bobsleigh athletes. The centerline of a bobsleigh track serves as a crucial dataset that defines the trajectory of the track. In contrast, the cross-sectional shapes are relatively standardized and can be generated using industrial software based on common templates bi:12 .
@@
This study focuses on typical bobsleigh track centerlines that determine the trajectory characteristics of the track and proposes a method for generating 3D centerlines from 2D centerlines. In accordance with international track design regulations, an optimization problem is formulated that considers factors such as total track length, height difference, slope, and continuity. A PGD is used to solve the problem. The generated 3D centerline is quantitatively compared with actual track data, and the influence of different track characteristics and algorithmic parameters on the generation results is investigated. The proposed method provides data support for enhancing the diversity of theoretical research and improving the generalizability of training outcomes for bobsleigh athletes.
@@ 2.1 Formulation of the Optimization Problem
According to the International Olympic Bobsleigh Competition Rules bi:14 , a bobsleigh track consists of a sequence of straight and curved sections. These two types of track segments exhibit distinct geometric characteristics, which necessitate different approaches for their analysis and generation. Therefore, it is essential to segment the track into straight and curved sections. Since the curvature of a straight segment is zero, curvature serves as a key variable for distinguishing between straight and curved portions.
@@
For bobsleigh tracks, slope and length are two critical characteristics. For a straight segment, the slope $g_{k}$ remains constant throughout. For a curved segment, when the distance $d$ between two points is extremely small, the segment between them can be approximated as a straight line. Therefore, if the coordinates of the previous point are $(x_{k-1},y_{k-1},z_{k-1})$, and the current point’s $x$ and $y$ coordinates are $(x_{k},y_{k})$, then the $z$-coordinate of the current point can be approximated as $z_{k}=z_{k}-d\cdot g_{k})$. The distance $d$ is calculated as follows:
@@
$J=\min_{g_{k}}\Bigg{[}$ $a\left(L-\sum_{k=2}^{n}d_{k}\right)+b\left(H-\sum_{k=2}^{n}\sqrt{(z_{k}-z_{k-1})^{2}}\right)$ (3)
$+\frac{c}{n}\sum_{k=1}^{n}(K_{pk}-K_{sk})+d\left(\frac{1}{j}\sum_{m=1}^{j}g_{m}-\bar{g}\right)\Bigg{]}$
[…]
In this equation, $J$ represents the cost function to be minimized, and $j$ is the number of track segments. The parameters $a$, $b$, $c$, and $d$ are the weighting factors corresponding to the total length difference, height difference, curvature difference between the 2D and 3D centerlines, and the average slope difference, respectively. $g_{k_{\text{max}}}$ and $g_{k_{\text{min}}}$ denote the maximum and minimum allowable slopes. $d_{k}$ is the 3D distance between point $k$ and point $k-1$, calculated as follows:
@@
where $u=[\dot{x},\dot{y},\dot{z}]$ is the first derivative vector of the coordinates, and $v=[\ddot{x},\ddot{y},\ddot{z}]$ is the second derivative vector of the coordinates.
To solve the bobsleigh track optimization problem established, a PGD bi:15 is used for parameter optimization. Compared to more complex methods such as genetic algorithms or simulated annealing, gradient descent offers faster convergence and requires fewer hyperparameters, making it more practical for repeated evaluations during track generation.
In this study, the height difference $h_{m}$ of each track segment is a high-dimensional parameter (with dimensionality corresponding to the number of track segments), which introduces significant nonlinearity to the optimization problem, making the exact computation. Therefore, the variation of the cost function is employed as a substitute for gradient computation. This approach does not rely on the explicit calculation of partial derivatives but instead approximates the gradient through a numerical difference method.
More specifically, the direction of the gradient is approximated by evaluating the change in the cost function value between two successive iterations, rather than explicitly computing all partial derivatives. The underlying concept is similar to the finite difference method; however, it places greater emphasis on variations in the cost function itself rather than on exact gradients. The corresponding computation is given as follows:
@@
where $\eta$ is the learning rate, which is set to $10^{-5}$.
If the scaling factor is considered, it should also be updated using gradient descent as follows:
@@
Compared to SGD(Stochastic Gradient Descent) bi:sgd1 ; bi:sgd2 ; bi:sgd3 ; bi:sgd4 , PGD can be regarded as projecting the height difference of each segment during optimization to ensure that the corresponding slope of each segment falls within the prescribed limits, which in turn satisfies the overall constraint conditions. Although the projection is not explicitly performed on each variable (e.g., the height difference of each segment) at every gradient descent step, it is effectively realized by adjusting the slope of each segment to meet the constraints.
@@
{$h_{m}&=g_{k_{\text{min}}}\times d_{m},\quad&&g_{m}\leq g_{k_{\text{min}}}\\ h_{m}&=g_{k_{\text{max}}}\times d_{m},\quad&&g_{m}\geq g_{k_{\text{max}}}\end{aligned}\right.$ (11)
@@ 2.2 Numerical Procedure
Based on the optimization equation formulated in accordance with the track design rules and the PGD, the computational procedure is illustrated in the following figures. First, the initial segmental height differences and the corresponding cost function value are calculated. Then, using a randomly assigned initial gradient, the segmental height differences are updated, and the $z$-coordinates of all points are recalculated based on these updated values.
Subsequently, the PGD algorithm enters an iterative loop. At each iteration, it first checks whether the slope of each track segment falls within the specified bounds. If any segment’s slope exceeds the upper or lower limits, the projection method is applied to map it back into the feasible range. Using the projected slope values, the segmental height differences are updated, followed by recalculating the $z$-coordinates of all points.
@@
If a scaling factor is included in the optimization, the $x$ and $y$ are scaled at the beginning of each iteration, and the scaling factor is updated at the end of each iteration. The optimization process is illustrated in Figure 2.
Figure 2: Calculation flow diagram.
3 Results Analysis: Model Verification
@@ 3.1 Track with Real 2D Plane Data
Since some studies include real 2D plane data for a limited number of tracks bi:6 , this data can be used for solving the optimization problem. The generated track data are then compared with the actual track data from the studies. In this section, the impact of the track’s ascent sections on the results is not considered. Therefore, only the data from the track’s descent sections are used for track generation and comparison analysis. In addition, the generation of the track centerline is a non-convex optimization problem, which may have multiple local optima. As a result, the algorithm may converge to different solutions in different runs. While the overall trend of the generated centerlines remains consistent, local variations may occur between runs. This characteristic enables the generation of diverse track designs, providing multiple alternative options for practical use and enhancing the flexibility and applicability of track design. Therefore, three random results are generated each time for comparing with the actual track centerline.
3.1.1 Whistler Track
The actual design geometric parameters of the Whistler track are shown in Table 1.
Table 1: Actual geometric parameters of the Whistler track
Total Length (m) Height Difference (m) Maximum Slope (%) Average Slope (%)
1278 149.3 20.4 11.6
For the Whistler track, Figures 3 and 4 show the comparison of the actual track centerline’s 2D curvature and the actual and generated track centerlines’ 3D curvatures as a function of track distance. From Figures 3 and 4, it can be observed that there is no significant difference between the 2D curvature of the actual track centerline and the 3D curvature of both the actual and generated tracks across the track distance. Only small differences are found near the extreme values. Therefore, the impact of this factor is not further analyzed in the results.
Figure 3: Comparison of the planar curvature and spatial curvature of the actual Whistler track centerline with track distance.
Figure 4: Comparison of the 2D curvature and 3D curvature of the actual bobsleigh track centerline with track distance.
Table 2 shows the comparison of the generated and actual Whistler bobsleigh track centerline geometric parameters. From Table 2, it can be observed that the generated track’s total length, height difference, and average slope are all close to those of the actual track, with maximum errors of 1.7%, 3.2%, and 4.1%, respectively. Additionally, the slope values do not exceed the defined limits.
@@ 3.1.1 Whistler Track
Table 2: Comparison of geometric parameters of the generated and actual Whistler track centerlines
Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
Generated track 1 1291.1 150.9 No 0.1208
Generated track 2 1290.2 144.5 No 0.1148
Generated track 3 1290.7 148.4 No 0.1177
Actual track 1278 149.3 No 0.1160
Figure 5 shows the comparison of the track height variation along the distance for the generated and actual Whistler track centerlines. From Figure 5, it can be seen that the three generated track centerlines closely follow the trend of the actual track centerline, with the track height continuously decreasing as the track distance increases. Specifically, for Generated Track 1, the height is lower than the actual track up to approximately 400 m, but higher beyond this point. This is because the average slope of each segment beyond 400 m is higher than that of the actual track, resulting in a larger height difference. For Generated Track 3, the height closely matches the actual track up to approximately 100 m, then falls below the actual track between 100 m and 300 m, and exceeds the actual track between 300 m and 900 m, before aligning with the actual track again beyond 900 m. This is because the randomly assigned slopes and height differences are close to those of the actual track. Generated Track 2 is slightly lower than the actual track between 0–200 m and beyond 1025 m, but closely matches the actual track between 200 m and 1025 m. This is due to the randomly assigned slopes and height differences being lower than those of the actual track.
Figure 5: Comparison of the track height variation along the distance for the generated and actual Whistler track centerlines.
Figure 6 shows the comparison of the track slope variation along the distance for the generated and actual Whistler track centerlines. From Figure 6, it can be seen that according to the study bi:6 , at track distances of approximately 100m, the model calculated in this paper results in a centrifugal force that exceeds the maximum allowable centrifugal force. Although the track slope does not exceed the design slope value of 0.204, the higher average slope, shorter total length, and larger height difference contribute to the centrifugal force exceeding the maximum allowable value. The different generated track slopes in this algorithm exhibit consistent trends along the track distance, with only minor variations in amplitude. All generated track slopes are within the design value range, demonstrating that the algorithm can provide a rich variety of track generation results, and that relevant parameters can be constrained to meet the track design requirements.
Figure 6: Comparison of the track slope variation along the distance for the generated and actual Whistler track centerlines.
3.1.2 Censana Track
Similar to the results of Whistler track, the actual geometric parameters of the Cesana track are shown in Table 3.
Table 3: Actual geometric parameters of the Cesana track
@@ 3.1.2 Censana Track
Total Length (m) Height Difference (m) Maximum Slope (%) Average Slope (%)
1359.2 124.5 18.3 9.1
Table 4 shows the comparison of geometric parameters of the generated and actual Cesana track centerlines. From Table 4, the simulated values of the generated track’s total length, height difference, and average slope are all close to those of the actual track, with maximum errors of 0.81%, 1.2%, and 2.6%, respectively. Moreover, all slope values remain within the prescribed limits.
Table 4: The comparison of geometric parameters of the generated and actual Cesana track centerlines
Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
Generated track 1 1360.3 124.0 No 0.0934
Generated track 2 1359.0 123.1 No 0.0926
Generated track 3 1359.0 123.0 No 0.0930
Actual track 1359.2 124.5 Exceeds by approximately 10m 0.0910
Figure 7 shows the comparison between the track height along the centerline of the generated and actual Cesana tracks as a function of distance. From Figure 7, the three centerlines generated by the proposed algorithm all follow a similar trend to that of the actual track, with the track height decreasing as the distance along the track increases.
For the generated Track 1, the height is close to that of the actual track before approximately 600 m, higher than the actual track between approximately 600 m and 1000 m, and lower than the actual track beyond approximately 1000 m. This is because the randomly assigned slope values in the segment between approximately 600 m and 1000 m are on average greater than those of the actual track. The result for generated Track 2 is similar to that of Track 1, with an even better match to the actual track before approximately 600 m. For generated Track 3, the result is similar to that of Track 1 before approximately 600 m, closely aligned with the actual track between approximately 600 m and 1000 m, but shows noticeable deviation after approximately 1000 m. This is due to the relatively smaller slope values assigned between approximately 600 m and 1000 m and the slightly lower overall height compared to the actual track.
@@
Figure 7: Comparison of track height variation along the distance for the generated and actual Cesana track centerlines.
Figure 8 shows the comparison of track gradient along the centerlines between the generated and actual Cesana track. As shown in Figure 8, although the maximum designed gradient value is 0.183 bi:6 , the actual value exceeds 0.2, which may lead to a centrifugal force exceeding the allowable limit of 5g during cornering 8 . According to the calculation model presented in the study bi:6 , the maximum centrifugal force is approximately 4.5g, which can be attributed to the moderate height difference and sufficiently long total track length. While a lower gradient is critical for safety, an overly small gradient may reduce the differentiation in athlete performance, resulting in smaller performance gaps. The gradients of the tracks generated by the proposed algorithm follow a similar trend along the track distance, with only local variations in magnitude. All gradient values remain within the design limits, demonstrating that the gradient profile of the generated centerlines meets the required specifications.
@@
Figure 8: Comparison of the track slope variation along the distance for the generated and actual Cesana track centerlines.
@@
In summary, the proposed optimization equation can generate track centerlines with diverse characteristics based on real 2D data, and the trend of the generated centerline’s characteristic parameters is consistent with that of the actual track. The generated centerlines effectively cover the spatial domain of the original track centerlines. Simulation results on two selected tracks show that the length, height difference, and average gradient of the generated tracks are close to those of the actual tracks, with maximum deviations of 1.7%, 3.2%, and 4.1%, respectively. Moreover, the gradients remain within the specified design limits, thereby validating the effectiveness of the proposed track centerline generation algorithm. The generation results are primarily influenced by parameters such as the randomly assigned segment gradients, height difference, distance, average gradient, and maximum gradient. In addition, the two real tracks analyzed herein highlight potential directions for further optimization.

### v2

# Projected gradient descent–based optimization of 3D bobsleigh track centerlines from 2D data for simulation

@@ Abstract
This paper proposes a method for generating a 3-dimensional (3D) track centerline based on 2D centerline data. Incorporating international track design regulations, the method formulates an optimization problem that considers total track length, height difference, slope constraints, and geometric continuity. A Projected Gradient Descent (PGD) algorithm is used to solve the optimization problem. Within the selected trajectories and parameter ranges in this study, the proposed algorithm can compute centerlines whose characteristic parameter trends are consistent with those of the actual track, based on either actual or scaled 2D data. Compared with the actual track, the maximum errors ranges for length, height difference, and average gradient are 1.1%, 16.4%, and 13.5%. Track sensitivity to segmentation and height-difference weighting varies with data type, and appropriate selection of segmentation and weighting parameters effectively reduces deviations and enhances the geometric accuracy of the generated centerlines.
@@
Bobsleigh; track centerline generation; international Olympic bobsleigh competition rules; projected gradient descent
articletype: ARTICLE TEMPLATE
@@ 1 Introduction
Since bobsleigh was formally recognized as a Winter Olympics event in 1924, it has attracted widespread attention from participating countries (bi:bu4). Bobsleighs are unpowered vehicles that rely entirely on the initial push from athletes and the conversion of gravitational potential energy into kinetic energy as they descend the track (bi:1; 3; 5). The competition is timed to an accuracy of milliseconds, and the final times often differ by only a few tenths of a second(4). To minimize energy loss, athletes typically avoid using the brake unless necessary before crossing the finish line. The dynamics of the bobsleigh and the driver’s control skill are key factors influencing performance (bi:bu1; roberts2013).
@@
However, due to restrictions from official governing bodies and the high cost of maintaining bobsleigh tracks(7), even countries with access to tracks often do not allow athletes to train on them regularly. Typically, limited test runs are permitted only shortly before official competitions (bi:2; rempfler2015; 4; 6). Consequently, simulation-based training and performance analysis have become vital tools for optimizing bobsleigh design and improving athlete technique. Researchers have developed virtual tracks, bobsleigh-blade/ice interaction models, multibody dynamics models of bobsleighs, and driver models. These simulation frameworks are increasingly being used to support bobsleigh structural design and help athletes devise optimal racing strategies (bi:2). Besides, there are limited studies on optimizing the centerlines of bobsleigh tracks.
@@
Early bobsleigh simulation research was significantly constrained by the limited computational capabilities of CPUs at the time, often relying on simplified point-mass models moving on smooth hyperbolic surfaces (bi:3; bi:4). The study (bi:4) developed a partial surface model of a bobsleigh track, treating the bobsleigh as a point mass to analyze its dynamic response under external forces. However, the authors also acknowledged that due to the lack of available data, virtual training environments differ significantly from the real settings, limiting the variety of technical strategies athletes can efficiently practice. Based on this point-mass framework and real data from the Lillehammer track in Norway, The study (bi:5) further designed an optimal controller to simulate bobsleigh driving behavior, aiding in athlete training. To better capture the complex nonlinear dynamics of bobsleigh motion on narrow, twisted, banked, and icy tracks, a geometric model is proposed (bi:6), which represents the tracks as a sequence of discrete points arranged along a 3D spatial curve. Moreover, a comprehensive simulation framework incorporating a blade–ice contact friction model is designed and validated on the Cesana and Whistler tracks, offering valuable guidance for track design and athlete training. The study (bi:8) further validates its agreement between the simulated and experimental results on the Cesana track. However, the study did not involve real-time simulation, and the applicability of the proposed model to other official tracks worldwide was not examined. Subsequently, the geometric model is further extended into a multibody dynamics model (bi:2), where more detailed dynamic characteristics are considered, resulting in better driving fidelity. Recent advances in bobsleigh simulation have focused on improving blade–ice friction modeling and virtual training environments (bi:bu3; bi:7). The study (bi:10) developed longitudinal and lateral friction models through experiments, FEM and nonlinear regression, validating their effectiveness in simulators to evaluate driving styles. Some studies further integrate steering and visual feedback or design immersive systems for training and tourism (bi:2; bi:3; bi:10; bi:11). The study (bi:11) introduced a VR hammock system to mimic sliding sensations, while the studies (bi:12; bi:13) proposed automated methods for generating 3D track surfaces.
@@
However, most of the aforementioned studies focus on the simulation of only one or two tracks, and the conclusions drawn lack generalizability. Many researchers have acknowledged the necessity of validating their proposed models against simulation results from other tracks in order to derive more universally applicable insights for athlete training. Currently, track data are primarily obtained through a combination of field measurements and industrial software modeling, which involves significant labor and material costs. Moreover, most countries do not publicly disclose the official data of their bobsleigh tracks. Therefore, acquiring a greater variety of available official track data worldwide has become an urgent need for enhancing both the diversity of research findings in simulation analysis and the general applicability of virtual training outcomes for bobsleigh athletes. The centerline of a bobsleigh track serves as a crucial dataset that defines a trajectory of a track. In contrast, the cross-sectional shapes are relatively standardized and can be generated using industrial software based on common templates (bi:12).
@@
This study focuses on typical bobsleigh track centerlines that determine the trajectory characteristics of the track and proposes a method for generating 3D centerlines from 2D centerlines. In accordance with international track design regulations, an optimization problem is formulated that considers factors such as total track length, height difference, slope, and continuity. And a PGD algorithm is used to solve the problem. The generated 3D centerline is quantitatively compared with actual track data, and the influence of different track characteristics and algorithmic parameters on the generation results is investigated. The proposed method provides data support for enhancing the diversity of theoretical research and improving the generalizability of training outcomes for bobsleigh athletes.
@@ 2.1 Formulation of the optimization problem
According to the International Olympic Bobsleigh Competition Rules (bi:14), a bobsleigh track consists of a sequence of straight and curved sections. These two types of track segments exhibit distinct geometric characteristics, which necessitate different approaches for their analysis and generation. Therefore, it is essential to segment the track into straight and curved sections. Since the curvature of a straight segment is zero, curvature serves as a key variable for distinguishing between straight and curved portions.
@@
For bobsleigh tracks, slope and length are two critical characteristics. For a straight segment, the slope $g_{k}$ remains constant throughout. For a curved segment, when the distance $d$ between two points is extremely small, the segment between them can be approximated as a straight line.
Therefore, if the coordinates of the previous point are $(x_{k-1},y_{k-1},z_{k-1})$, and the current point’s $x$ and $y$ coordinates are $(x_{k},y_{k})$, then the $z$-coordinate of the current point can be approximated as $z_{k}=z_{k-1}-d\cdot g_{k}$. The distance $d$ is calculated as follows:
@@
$J=\min_{g_{k}}\Bigg[$ $a\left(L-\sum_{k=2}^{n}d_{k}\right)^{2}+b\left(H-\sum_{k=2}^{n}\sqrt{(z_{k}-z_{k-1})^{2}}\right)^{2}$ (3)
$+c\left(\frac{1}{n}\sum_{k=1}^{n}(K_{pk}-K_{sk})\right)^{2}+d\left(\frac{1}{j}\sum_{m=1}^{j}g_{m}-\bar{g}\right)^{2}\Bigg]$
$\text{s.t.}\quad g_{k_{\text{min}}}\leq g_{k}\leq g_{k_{\text{max}}}$
In this equation, $J$ represents the cost function to be minimized, and $j$ is the number of track segments. The parameters $a$, $b$, $c$, and $d$ are the weighting factors corresponding to the total length difference, height difference, curvature difference between the 2D and 3D centerlines, and the average slope difference, respectively, $g_{k_{\text{max}}}$ and $g_{k_{\text{min}}}$ denote the maximum and minimum allowable slopes, $d_{k}$ is the 3D distance between point $k$ and point k1k-1, calculated as follows:
@@
where $u=[\dot{x},\dot{y},\dot{z}]$ is the first derivative vector of the coordinates, and $v=[\ddot{x},\ddot{y},\ddot{z}]$ is the second derivative vector of the coordinates. To solve the bobsleigh track optimization problem established, a PGD algorithm (9; 10) is used for parameter optimization. Compared to more complex methods such as genetic algorithms or simulated annealing, gradient descent offers faster convergence and requires fewer hyperparameters, making it more practical for repeated evaluations during track generation.
In this study, the height difference $h_{m}$ of each track segment is a high-dimensional parameter (with dimensionality corresponding to the number of track segments), which introduces significant nonlinearity to the optimization problem, making the exact computation. Therefore, the variation of the cost function is employed as a substitute for gradient computation. This approach does not rely on the explicit calculation of partial derivatives but instead approximates the gradient through a numerical difference method. The corresponding computation is given as follows:
@@
where $\eta$ is the learning rate, which is set to $10^{-5}$. If the scaling factor is considered, it should also be updated using gradient descent as follows:
@@
Compared to SGD (Stochastic Gradient Descent) algorithms (11; 12; 13), PGD algorithms can be regarded as projecting the height difference of each segment during optimization to ensure that the corresponding slope of each segment falls within the prescribed limits, which in turn satisfies the overall constraint conditions. Although the projection is not explicitly performed on each variable (e.g., the height difference of each segment) at every gradient descent step, it is effectively realized by adjusting the slope of each segment to meet the constraints.
@@
$\left\{\begin{aligned} h_{m}&=g_{k_{\text{min}}}\times d_{m},\quad&&g_{m}\leq g_{k_{\text{min}}}\\ h_{m}&=g_{k_{\text{max}}}\times d_{m},\quad&&g_{m}\geq g_{k_{\text{max}}}\end{aligned}\right$. (12)
@@ 2.2 Numerical Procedure
Based on the optimization equation formulated in accordance with the track design rules and the PGD algorithm, the computational procedure is illustrated in figure 2. Firstly, the initial segmental height differences and the corresponding cost function value are calculated. Then, using a assigned initial gradient to update segmental height differences, and the $z$-coordinates of all points are recalculated based on these updated values.
Subsequently, the program enters the PGD algorithm iterative loop. At each iteration, it first checks whether the slope of each track segment falls within the specified bounds. If any segment’s slope exceeds the upper or lower limits, the projection method is applied to map it back into the feasible range. Using the projected slope values, to update the segmental height differences, followed by recalculating the $z$-coordinates of all points.
@@
As shown in the dashed box in the figure, if a scaling factor is included in the optimization, the $x$ and $y$ are scaled at the beginning of each iteration, and the scaling factor is updated at the end of each iteration.
Figure 2: Numerical procedure.
3 Model Verification
@@ 3.1 Track with Real 2D Plane Data
Since some studies include real 2D plane data for a limited number of tracks (bi:6), this data can be used for solving the optimization problem. The generated track data are then compared with the actual track data from the studies. Due to the fact that the actual track height difference parameter that can be obtained can only be reflected in the descending section of the track, and the descending section of the track is the key factor affecting the competition results. Therefore, in this section, the impact of the track’s final ascent sections on the results is not considered.
Table 1 present comparisons of the geometric parameters of the generated and actual bobsleigh track centerlines for the Whistler and Cesana tracks. As shown in table 1, for the Whistler track, the generated tracks demonstrate strong consistency with the actual tracks in terms of total length, height difference, and average slope, with maximum deviations of 0.7%, 0.4%, and 0.5%, respectively. For the Cesana track, the corresponding deviations are 0.2%, 0.6%, and 1.1%. Furthermore, the generated tracks exhibit smaller average differences between the 2D and 3D curvature compared with their actual counterparts, while the slope values remain within the defined design limits.
@@
Table 1: Comparison of geometric parameters of the generated and actual Whistler track.
@@
Total Length (m) Height Difference (m) Average difference between 2D and 3D curvature Max Slope Average Slope
Generated Track (Whistler) 1286.7 148.7 3.97×10-4 0.152 0.1182
Actual Track (Whistler) 1278 149.3 9.76×10-4 0.204 0.1176
Generated Track (Cesana) 1356.7 123.7 2.52×10-4 0.147 0.0920
Actual Track (Cesana) 1359.2 124.5 5.04×10-4 0.183 0.0910
@@
Figure 3 shows the comparison of the track height variation and slope along the distance for the generated and actual track centerlines. Among them, Figure 3(a) is the Whistler track, and Figure 3(b) is the Cesana track.
(a) Whistler track.
(b) Cesana track.
@@
Figure 3: The comparison of the track height variation and slope along the distance for the generated and actual track centerlines.
For the generated Whistler track, the height is lower than that of the actual track from 0 to approximately 245m, but higher between 245m and 1140m. Beyond 1140m, the heights converge. This is mainly attributed to the influence of slope variations. From the overall slope variation curve, the slopes of the generated track remain within the reasonable design range. The slopes of the actual and generated tracks alternate, resulting in height differences. For example, between 0 and 205m, the slope of the generated track is smaller than that of the actual track, leading to a lower height in this section. In contrast, from 205m to 245m, the slope of the generated track exceeds that of the actual track, causing the height to gradually approach that of the actual track.
For the generated Cesana track, the height is close to that of the actual track from 0 to approximately 115m, but higher between 115m and 1266m. Beyond 1266m, the heights converge as well. In addition, due to slope variations, the slope of the generated track is greater than that of the actual track in most regions between 115m and 950m, resulting in higher track heights. In contrast, between 950m and 1110m, the slope of the generated track is smaller than that of the actual track, leading to a gradual convergence of height toward that of the actual track.
According to the study by bi:6, for the Whistler track, when the track distance is approximately 100m, the centrifugal force of the actual track exceeds the maximum allowable value. Although the track slope does not exceed the design limit of 0.204, a higher average slope, shorter total length, and larger height difference contribute to the centrifugal force surpassing the permissible threshold of 5g. In the case of the Cesana track, although its maximum design slope is 0.183, the actual slope exceeds 0.2, which may cause the centrifugal force during turns to exceed the allowable limit. The maximum centrifugal force of the actual track is about 4.5g, which remains within a reasonable range owing to the appropriate height difference and sufficient total track length. Although a smaller slope is critical for safety, an excessively small slope may reduce performance differences among athletes, thereby narrowing performance gaps.
Figures 4 shows the comparison of trajectory differences between the generated and actual track centerlines, in terms of 2D and 3D curvature variation along the track distance. Among them, Figure 4(a) is the Whistler track, and Figure 4(b) is the Cesana track. As shown in the figures, in most regions, the differences between the 2D and 3D curvature of the generated track are smaller than or close to those of the actual track. This prevents excessive fluctuations in the slope in the z-direction and mitigates large vertical accelerations, thereby improving the safety and smoothness of the track design.
(a) Whistler track.
(b) Cesana track.
Figure 4: The comparison of trajectory differences between the generated and actual track centerlines.
@@
In summary, the proposed optimization equation can generate track centerlines based on real 2D data, and the proposed algorithm can compute centerlines whose characteristic parameter trends are consistent with those of the actual track. Simulation results on two selected tracks show that the length, height difference, and average gradient of the generated tracks are close to those of the actual tracks. Moreover, the slopes remain within the specified design limits, thereby validating the effectiveness of the proposed track centerline generation algorithm. The generation results are primarily influenced by parameters such as height difference, distance, average gradient, maximum gradient, average differences between the 2D and 3D curvature. In addition, the two real tracks analyzed herein highlight potential directions for further optimization.